in the fu tbove the end -167 and and What is All play the action that the point the bal wher
Answer #1

Given the position of the object t seconds after the launch (which is also its height above the ground) is given by, s(t)= -1Graph of the function s(t)= -16° +36 +3 as shown below: Y 40+ $(t) = -16r2 + 367 +3 (b) Find the height of the ball when it r(-16° + 36+ + 3) - (-168°) + (361) + . (3) =-32t +36 + 0 = -32t +36 Then, -32t +36 = 0 32t = 36 Substitute t = - sec in s(t)=+ 81+12 + = 23.25 ft Hence, the height of the ball when it reaches the high point is 23.25 ft.(c) To find how long does the catcher have to wait to catch the ball. Put s(t)=0. -16t2 + 36t +3 = 0 -360 (36)2 - 4.(-16)(3)